OpenStudy (anonymous):

Write an equation in standard form using only integers for a line that passes through the point (4,3) and is parallel to the line: -10x -2y=-5 Enter A and C where your answer is Ax+y=C

OpenStudy (imstuck):

First thing to do is to find the slope of the line they gave you. Once you know that, you know the slope of the line you are looking for, cuz parallel lines have the same slope.

OpenStudy (anonymous):

so i got -5 as the slope

OpenStudy (imstuck):

right. Now use the point they gave you to write an equation in point-slope form. Can you do that? This problem takes several steps to get to the answer in the form they want it in.

OpenStudy (imstuck):

What is the equation in point-slope form?

OpenStudy (anonymous):

oh wait i kow how to do this

OpenStudy (imstuck):

I bet you do! I'll wait for your equation!

OpenStudy (anonymous):

yeah the y=mx+b

OpenStudy (imstuck):

you have to use the other one first. That one is slope-intercept, but you need to find it in point-slope first using the point they gave you.

OpenStudy (imstuck):

First things first.

OpenStudy (anonymous):

oh alright, but wait that's not the question I'm having trouble with, i just answered this one. Do you think you could help me with this one : Write an equation in standard form using only integers for a line that passes through the point (7,6) and is perpendicular to the line: 7x +2y=-4 Enter B and C where your answer is -2x+By=C

OpenStudy (imstuck):

sure! find the slope of this one first. What is the slope of the line perpendicular to the one they gave you?

OpenStudy (anonymous):

-7/2?

OpenStudy (imstuck):

Thats the slope of the line they gave you? What about the slope of the line that is perpendicular to that one?

OpenStudy (imstuck):

Do you know the rule for perpendicular lines andd their slopes?

OpenStudy (anonymous):

no :/ that's why i am having trouble. i'm not sure how to solve it

OpenStudy (imstuck):

Lines that are perpendicular to one another have slopes that are opposite reciprocals. Meaning if the slope of the line they gave you is -7/2, the line perpendicular to that line has a slope of +2/7. That's the rule.

OpenStudy (imstuck):

It holds true for all lines that are perp. to one another.

OpenStudy (anonymous):

alright so what do i do from there?

OpenStudy (imstuck):

Now that we found the slope of the line perpendicular to the one they gave you, use the point in the point-slope formula and write the equation using that new slope.

OpenStudy (anonymous):

how does that look like?

OpenStudy (imstuck):

Here, like this: